OpenStudy (jadareese_18):

Write the exponential function f(x)=-3 x 4^(1-x) in the form f(x)=ab^x. (In the exponential form that the function is written in above, the x between the -3 and 4 is a multiplication symbol)

jhonyy9 (jhonyy9):

(1-x) is the exponent of 4 ? or is the power of x too same with this 4 ? or there are -3 time 4^(1-x) ?

jhonyy9 (jhonyy9):

f(x)=a*b^x where -3 mean this a and 4 mean b and (1-x) mean this exponent of x ?

OpenStudy (jadareese_18):

Does this help? @jhonyy9

jhonyy9 (jhonyy9):

yes ty. this is more clearly now so -3 is right but there is 4 on exponent of (1-x) for this case of 4 use this property of exponents a^(x-y) = (a^x)/(a^y) hope helped

OpenStudy (3mar):